A dehydrated patient needs a 7.9% saline IV. Unfortunately, the hospital only has bags of 3% and 10% saline solutions. How many liters of each of these solutions should be mixed together to yield 3 liters of the desired concentration?
Answer by jorel1380(3719)

3x.079=.237
x+y=3
3x+10y=23.7
3x+3y=9
7y=14.7
y=2.1
x=.9
You need .9 liters of 3% solution and 2.1 liters of 10% solution to make 3 liters of 7.9% solution..